### v1

# Using diffusion model as constraint: Empower Image Restoration Network Training with Diffusion Model

Abstract
Image restoration has made marvelous progress with the advent of deep learning. Previous methods usually rely on designing powerful network architecture to elevate performance, however, the natural visual effect of the restored results is limited by color and texture distortions. Besides the visual perceptual quality, the semantic perception recovery is an important but often overlooked perspective of restored image, which is crucial for the deployment in high-level tasks. In this paper, we propose a new perspective to resort these issues by introducing a naturalness-oriented and semantic-aware optimization mechanism, dubbed DiffLoss. Specifically, inspired by the powerful distribution coverage capability of the diffusion model for natural image generation, we exploit the Markov chain sampling property of diffusion model and project the restored results of existing networks into the sampling space. Besides, we reveal that the bottleneck feature of diffusion models, also dubbed h-space feature, is a natural high-level semantic space. We delve into this property and propose a semantic-aware loss to further unlock its potential of semantic perception recovery, which paves the way to connect image restoration task and downstream high-level recognition task. With these two strategies, the DiffLoss can endow existing restoration methods with both more natural and semantic-aware results. We verify the effectiveness of our method on substantial common image restoration tasks and benchmarks. Code will be available at https://github.com/JosephTiTan/DiffLoss.
Keywords:
Image restoration Diffusion model Perception quality Low-for-high
1 Introduction
In complex imaging environments, the quality of imaging often suffers from unpredictable degradations, such as low-light conditions, heterogeneous media, leading to information loss both in content and color and unnatural visual effects. Besides, these degradations also causes adverse effects on high-level computer vision tasks such as object detection [8, 32] and scene understanding [53, 52]. Thus, it is indispensable to develop effective algorithms to recover high-quality degradation-free images in natural appearance, which is referred to as image restoration. This task has been a longstanding and challenging problem, drawing the attention of researchers.
In fact, estimating the clean image from the degraded input is an ill-posed and challenging problem. Previous traditional methods utilized the inherent feature information of degraded images to design algorithms with prior knowledge for image restoration[4, 23, 16, 15]. However, these hand-crafted image priors are drawn from specific observations with limited robustness, which may not be reliable for modeling the intrinsic characteristics of degraded images. With the advent of deep learning, image restoration has made marvelous progress, where most of them rely on designing advanced architectures to learn degradation to clean mappings [31, 38, 14, 64, 21, 67]. However, due to the severe degradation, the naturalness of the restored results is limited with color and texture distortions, as shown in Fig. 1. This is challenging, if not impossible, to solve in the current framework, wherein we employ mapping-based deep learning paradigm and limited area data. Besides, due to the common practice of relying solely on visual quality metrics during the restoration process, the semantic perception recovery is usually absent in the recovered image, resulting in poor performance in downstream high-level tasks even with seemingly high visual quality.
Figure 1: Visual comparison between baselines with and without our DiffLoss. (a) is produced with IAT[11] on LOL dataset[61]; (b) is produced with MSBDN[14] on NH-HAZE dataset [2]. Without our DiffLoss, the visual result suffers from unnaturalness problem, with color shift and content artifacts. DiffLoss largely alleviates this issue with more natural results. Need to mention that DiffLoss is an optimization strategy. The improvement should be compared with the baseline method, instead of other restoration methods or the ground-truth images.
To this end, we argue that, to meet the challenging information loss issue caused by severe degradation and limited data and model capacity, image restoration methods should be explicitly empowered with models that has strong prior on natural and clean information distribution modeling. Inspiringly, the emerged diffusion model has exhibited strong natural image generation capability [13] and possible semantic potential [3], which open the avenue to solve these problems.
Figure 2: Illustration of images after modifying h-space features with SVD. As h-space changes, the semantic information of the image transitions from an axolotl to a cat and gradually loses its original semantics completely. DiffLoss preserves the original semantic information of the image through h-space.
In this paper, we propose a new perspective to resort these issues by introducing a naturalness-oriented and semantic-aware optimization mechanism, dubbed DiffLoss. Specifically, drawing inspiration from the diffusion model’s remarkable capacity to cover distributions in natural image generation, we leverage the Markov chain sampling characteristic of the diffusion model to project the restored results of existing networks into the sampling space. Besides, we reveal that the bottleneck feature of diffusion models, also dubbed h-space feature, is a natural high-level semantic space, as shown in Fig. 2. We investigate this property and introduce a semantic-aware loss to further unleash its potential in recovering semantic perception. This breakthrough paves the way for bridging the gap between image restoration tasks and downstream high-level recognition tasks. With these two strategies, the DiffLoss can endow existing restoration methods with both more natural and semantic-aware results.
Different from previous methods that focus on dedicated model design or directly employ diffusion model as the restoration model, our DiffLoss works as a general and novel auxiliary training mechanism, to boost the ability of existing methods to achieve better visibility effects and preservation of semantic information. Specifically, we employ the unconditional diffusion model pre-trained on ImageNet dataset as optimization prior to exploit its immense clean data distribution. The diffusion model takes the restored image and ground truth as inputs and projects them into the distribution sampling space and semantic h-space. Within these spaces, we can get more natural result and preserve more semantic information by pulling the intermediate sampling distribution and h-space of the restored image closer to that of the corresponding clear image.
Besides, it’s especially beneficial for empowering parameter-limited models as it involves naturalness prior. Our method also involves no additional computations in the inference stage. We verify the effectiveness of our method on substantial common image restoration tasks, including image dehazing, image deraining, and low-light image enhancement.
Overall, our contributions can be summarized as follows:
  • We introduce the naturalness and recognition-aware modeling paradigm and into restoration network to alleviate the unnaturalness problem and recognition-unfriendly issue of existing methods by embedding the diffusion model as auxiliary training mechanism.
  • Specifically, our approach leverages the fixed diffusion model to enable tractable sampling and the extraction of semantic information, and it yields a natural sample and recognition-aware restoration when the DiffLoss is minimized.
  • Extensive experiments demonstrate that our DiffLoss empowers existing restoration methods and helps training efficient models and improving the Classification ability on data with varying of degradation without involving additional computations in the inference stage.
2 Related work
2.1 Image Restoration
Image restoration refers to the process of improving the quality of a degraded or damaged image by removing various types of degradations. With the advent of deep learning, numerous methods leveraging deep neural networks for image restoration have emerged, such as low-light enhancement[60, 69, 70, 73], super-resolution[30, 37, 71, 72], dehazing[4, 7, 15, 23, 33, 31, 68, 66], deraining[19, 35, 40, 45, 48], deblurring[5, 9, 20, 29, 49]. However, the severe degradation limits the naturalness of restored results, causing color and texture distortions. Moreover, relying solely on visual quality metrics during the restoration process, without considering the semantic aspects of quality, has negatively impacted performance in downstream tasks.
2.2 Diffusion models for image restoration
Existing diffusion models used in image restoration tasks can be roughly categorized into two classes: conditional diffusion and unconditional diffusion.
Conditional DDPMs [10, 51, 63, 50, 43, 27, 62] are usually combined with specific image restoration task, such as image super-resolution [51], image deblurring [63], and image deraining [62]. SR3 trains a conditional diffusion model for image super-resolution with the low-resolution images as condition. Whang et al. [63] proposed the "predict and refine" strategy and learned the residual with conditional diffusion model in image deblurring task. RainDiffusion [62] combines cycle-consistent architecture with diffusion model to achieve unsupervised image deraining.
Unconditional DDPMs [39, 27, 58, 55, 41] are usually integrated into general image restoration task. For example, RePaint [39] solves inpainting problem by employing unconditional diffusion process in the unmasked region and reverse back to solve boundary inconsistency. DDRM [27] uses SVD to decompose the degradation operators and embeds unconditional diffusion model into unsupervised posterior sampling method to solve various linear inverse problems. DDNM [58] applies range-null space decomposition to degraded images and refines only the null-space contents during the reverse process to yield diverse results.
However, all these methods attempt to integrate diffusion models into the network architecture that has the limit of slow inference speed and accounting huge memory cost when deploying. We circumvent these limitations from a new perspective by exploring diffusion model as an auxiliary training mechanism to empower the learning capability of existing image restoration networks, without involving additional computations in the inference stage.
3 Motivation
Inspiration from 3D generation. Our DiffLoss is similar to the Score Distillation Sampling (SDS) in Dreamfusion  [44] that perform text-to-3D synthesis with NeRF model. By using a pretrained 2D text-to-image diffusion model to produce a prior optimization loss and generate corresponding 3D object given a text, the NeRF model distills prior information of natural images from the diffusion model. Similarly, we leverage the unconditional diffusion model pretrained on large-scale ImageNet dataset as optimization prior. Different from SDS, our method is exploited in the 2D image restoration tasks instead of 3D generation and our method is a supervised regression task with degraded image as input and clear image as supervision, while Dreamfusion is a zero-shot generation task with only text prompt as input.
Why DiffLoss can empower restoration networks. Previous loss [12, 17] that assist restoration networks have the following drawbacks: (1) L1 or L2 loss works in pixel space, which may produce images deviated from natural distribution. (2) The VGG16 used in perceptual loss is pre-trained for high-level task, instead of low-level image restoration task. (3) Adversarial loss treats restoration network as generator and inserts an additional discriminator network and needs to train a discriminator for every restoration dataset, which is troublesome and time-consuming.
Thanks to the strong generation ability of DDPMs and the immense data samples in the training dataset, the unconditional diffusion model pre-trained on ImageNet dataset possesses world-wide natural image data distribution and has been applied in image restoration tasks  [39, 27, 58, 55, 41]. The wide distribution and mode convergence property enables it to be a powerful and general image prior, and fits for both general and specific low-level image restoration tasks. Besides, h-space in diffusion model is also found to reflect the semantics of images. Hence, by pulling the intermediate sampling stages and h-space closer to that of clear images and leveraging the distribution sampling property of diffusion model, the restored results can be optimized to be more natural and recognition-aware, which is difficult to achieve for the previous methods. In contrast to previous usage of unconditional diffusion model in low-level vision usually directly exploits its generation ability by incorporating diffusion model into the network backbone, our method propose the "implicit" usage of unconditional diffusion model and employ it as auxiliary training mechanism to solve the unnaturalness problem in image restoration. More analysis is provided in the supplementary materials.
Figure 3: Overview structure of our method. For any existing restoration network, we train it with the aid of our DiffLoss to achieve higher performance and more natural visual results. The parameter of DiffLoss is frozen during training stage. During inference stage, we only have the optimized restoration network, without involving the DiffLoss.
4 Methodology
In this section, we first briefly introduce the denoising diffusion probabilistic models, followed by a detailed presentation of the DiffLoss we propose.
4.1 Denoising Diffusion Probabilistic Models
DDPM is a latent variable model specified by a T-step Markov chain, which approximates a data distribution $q(x)$ with a model $f_{\theta}(\cdot)$. It contains two processes: the forward diffusion process and the reverse denoise process.
The forward diffusion process. The forward diffusion process starts from a clean data sample $x_{0}$ and repeatedly injects Gaussian noise according to the transition kernel $q(x_{t}|x_{t-1})$ as follows:
$\small q(x_{t}|x_{t-1})=N(x_{t};\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})I),\$ (1)
where $\alpha_{t}$ can be learned by reparameterization [28] or held constant as hyper-parameters, controlling the variance of noise added at each step. From the Gaussian diffusion process, we can derivate closed-form expressions for the marginal distribution $q(x_{t}|x_{0})$ and the reverse diffusion step $q(x_{t-1}|x_{t},x_{0})$ as follows:
$q(x_{t}|x_{0})=N(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I),$ (2)
$q(x_{t-1}|x_{t},x_{0})=N(x_{t-1};\tilde{\mu}_{t}(x_{t},x_{0}),\tilde{\beta}_{t}I),$ (3)
where $\tilde{\mu}_{t}(x_{t},x_{0}):=\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}x_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x_{t}$, $\tilde{\beta}_{t}:=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}(1-\alpha_{t})$, and $\bar{\alpha_{t}}:={\textstyle\prod_{s=1}^{t}}\alpha_{s}$.
Note that the above-defined forward diffusion formulation has no learnable parameters, and the reverse diffusion step cannot be applied due to having no access to $x_{0}$ in the inference stage. Therefore, we further introduce the learnable reverse denoise process for estimating $x_{0}$ from $x_{T}$.
The reverse denoise process. The DDPM is trained to reverse the process in Equation 1 by learning the denoise network $f_{\theta}$ in the reverse process. Specifically, the denoise network estimates $f_{\theta}(x_{t},t)$ to replace $x_{0}$ in Equation 3. Note that $f_{\theta}(x_{t},t)$ directly predicts the Gaussian noise $\varepsilon$, instead of $x_{0}$. While, the estimated $\varepsilon$ deterministicly corresponds to $\hat{x}_{0}$ via Equation 2.
$p_{\theta}(x_{t-1}|x_{t})$ $=q(x_{t-1}|x_{t},f_{\theta}(x_{t},t))$ (4)
$=N(x_{t-1};\mu_{\theta}(x_{t},t),{\textstyle\sum_{\theta}}(x_{t},t)).$
$\small\mu_{\theta}(x_{t},t)=\tilde{\mu_{t}}(x_{t},x_{0}),{\textstyle\sum_{\theta}}(x_{t},t)=\tilde{\beta_{t}}I.$ (5)
Similarly, the mean and variance in the reverse Gaussian distribution 4 can be determined by replacing $x_{0}$ in $\tilde{\mu}_{t}(x_{t},x_{0})$ and $\tilde{\beta}_{t}$ with the learned $\hat{x}_{0}$
Training objective and sampling process. As mentioned above, $f_{\theta}(x_{t},t)$ is trained to approach the Gaussian noise $\varepsilon$. Thus the final training objective is:
$L=E_{t,x_{0},\varepsilon}\left\|\varepsilon-f_{\theta}(x_{t},t)\right\|_{1}.$ (6)
The sampling process in the inference stage is done by running the reverse process. Starting from a pure Gaussian noise $x_{T}$, we iteratively apply the reverse denoise transition $p_{\theta}(x_{t-1}|x_{t})$ T times, and finally get the clear output $x_{0}$.
Figure 4: Detailed design of DiffLoss. We devise our DiffLoss with $t$ step forward process and one step reverse process. The $t$ step forward process can be directly achieved with Eq. 2. After projecting the restored image and the clear image into intermediate noisy images with Eq. 2, these noisy images are further fed into the denoising UNet. The DiffLoss is designed to pull the output of the denoising UNet and the bottleneck feature closer.
4.2 Overview structure
As presented in Fig. 3, in our main setting, given the degraded image $y$, its corresponding clear image $x$ , any existing restoration network $g_{\phi}(\cdot)$, and pre-trained diffusion model $f_{\theta}(\cdot)$ (with fixed parameter $\theta$), we first get the restored output $z=g_{\phi}(y)$ from the restoration network. Then DiffLoss works as auxiliary training mechanism providing naturalness modeling ability and for restoration networks.
4.3 Detailed design of DiffLoss
Originally, the forward diffusion process translates a clean data sample $x_{0}$ into Gaussian noise $\varepsilon$ by gradually adding Gaussian noise to $x_{0}$ in a parameter-free manner. The reverse denoise process is trained to sample from Gaussian noise $\varepsilon$ to generate clean images via gradually removing noise with the denoise network $f_{\theta}$. However, both of these two processes have asymmetric input-output pairs and thousands of iterative steps. These properties are not suitable for direct loss design. For example, direct applying the reverse denoise process $T$ times is time-consuming and impacts the backpropagation of gradients. Besides, since the DiffLoss takes the restored result $z$ as input, it should start from forward process, and connect to the reverse process in a proper way. Because single forward process is parameter-free and can be used as a learnable loss.
To this end, we redesign diffusion model delicately. We employ $t$ step forward diffusion process and one step reverse denoise step to minimize time-consuming as well as get symmetric image-image input-output pairs. Specifically, as shown in Fig. 3 and Fig. 4, we integrate the forward diffusion process and reverse denoise process. In the forward diffusion process, we get the intermediate noisy image $x_{t}$ via $q(x_{t}|x_{0},\varepsilon)$, which is expressed as Equation 7.
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{(1-\bar{\alpha}_{t})}\varepsilon.$ (7)
$x_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}x_{t}-\sqrt{(\frac{1}{\bar{\alpha}_{t}}-1)}\varepsilon.$ (8)
Then comes the reverse denoise process. Note that Eq. 8 provides the way of reconstructing $x_{0}$ back. However, we need to implement Eq. 8 in a learnable way. Accordingly, we devise two ways by respectively replacing $x_{t}$ and $\varepsilon$ with the diffusion model learned ones, as shown in Fig. 3.
In the first way, starting form $x_{t}$, we get the pseudo Gaussian noise $\hat{\varepsilon}=f_{\theta}(x_{t},t)$. Then, the pseudo Gaussian noise map $\hat{\varepsilon}$ is employed to replace $\varepsilon$ in Eq. 8. The first inverse way $q(\hat{x_{0}}|x_{t},\hat{\varepsilon})$ is expressed as follows:
$\hat{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}x_{t}-\sqrt{(\frac{1}{\bar{\alpha}_{t}}-1)}\hat{\varepsilon}.$ (9)
The second manner starts from $x_{t}$ to $\hat{x}_{t-1}$ via $q(\hat{x}_{t-1}|x_{t},f_{\theta}(x_{t},t))$, and finally get the output $\hat{x_{0}}$ from $\hat{x}_{t-1}$ with Eq. 10:
$\hat{\hat{x}}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t-1}}}\hat{x}_{t-1}-\sqrt{(\frac{1}{\bar{\alpha}_{t-1}}-1)}\varepsilon,$ (10)
Figure 5: Visualization of the intermediate and output images of DiffLoss at different step t. Despite input a clear image into DiffLoss module, the intermediate stage reconstructed $\hat{x}_{0}$ and $\hat{\hat{x}}_{0}$ are not complete reconstructions of the input image.
Besides, we show $x_{t}$, $\hat{x}_{t-1}$, $\hat{x}_{0}$, and $\hat{\hat{x}}_{0}$ in Fig. 5 to visualize how the intermediate sampling stage changes with $t$. Note that although the visual difference between $x_{t}$ and $\hat{x}_{t}$ is negligible, the reconstructed $\hat{x}_{0}$ and $\hat{\hat{x}}_{0}$ are obviously different, and are not complete reconstruction of $x$. We can also direct constrain on $\hat{\varepsilon}$ and $\hat{x}_{t-1}$ without reconstructing back to $\hat{x}_{0}$ and $\hat{\hat{x}}_{0}$, which are work equally. In this work, we choose current way to better visually illustrate the difference of the intermediate sampling stages. In this way, the restoration networks can harness the generative capabilities of the diffusion model, as well as the sampling space’s reflection of the natural attributes of images, to obtain restored results with more naturalness.
If we decompose $f_{\theta}(x_{t},t)$ into an encoder $\mathcal{E}_{\theta}(\cdot)$ and a decoder $\mathcal{D}_{\theta}(\cdot)$ , then after obtaining the noisy version of clear image $\hat{x}_{t}$ and restored image $\hat{z}_{t}$, the form of h-space feature of them can be written as the following equations respectively:
$\hat{h}_{clr}^{t}=\mathcal{E}_{\theta}(x_{t},t),$ (11)
$\hat{h}_{rst}^{t}=\mathcal{E}_{\theta}(z_{t},t),$ (12)
where $\hat{h}_{clr}^{t}$ and $\hat{h}_{rst}^{t}$ reflect the semantic information of the clear image and restored image, respectively. By reducing the gap between these two terms, we can preserve more semantic information in the restored image.
[…]
$L_{DiffLoss}$ $=\left\|\hat{x}_{0}-\hat{z}_{0}\right\|_{2}+\left\|\hat{\hat{x}}_{0}-\hat{\hat{z}}_{0}\right\|_{2}+\lambda\left\|\hat{h}_{clr}^{t}-\hat{h}_{rst}^{t}\right\|_{2},$ (13)
$L_{total}$ $=\left\|x-z\right\|_{2}+\gamma L_{DiffLoss}.$ (14)
where $\lambda$ and $\gamma$ are weight factors. We empirically set the weight $\lambda=0.01$ and $\gamma=0.001$. Besides, we also try adaptive weight which is correlated with timestep $t$. Both strategies perform similarly. A detailed ablation on the weight of DiffLoss is conducted in the experimental part. By employing these two strategies, the DiffLoss has the potential to enhance existing restoration methods, yielding restored images that are not only more natural but also reserve more semantic information.
5 Experiments
In this section, we first introduce the datasets and implement details of our experiment. Then, we conduct experiments on existing restoration methods, including low light enhancement, image deraining and image dehazing, with and without our DiffLoss. To validate the effectiveness of our approach in low-level tasks, we conducted tests on the efficiency model. Additionally, we examined the performance of the model employing our method in the image classification task with degraded data. Experiments on several baselines and benchmarks demonstrate the effectiveness of our DiffLoss.
Table 1: Quantitative comparison on three datasets. $\uparrow$ indicates that the larger the value, the better the performance. ``"``"``*"` ` ∗ " refers to the efficient model of the task.
Dehazing Deraining Low light enhancement
Method Dense-Haze Method Rain100H Method LOL
[…]
AOD-Net[31] 13.14 0.4144 DerainNet[18] 14.92 0.5920 EnlightenGAN[26] 17.48 0.6510
FFA-Net[46] 14.39 0.4524 RESCAN[35] 26.36 0.7860 SwinIR[36] 17.81 0.7230
AECR-Net[64] 15.80 0.4660 PreNet[48] 26.77 0.8580 AirNet[34] 18.18 0.7350
FSDGN[67] 14.34 0.4010 RCD-Net[57] 17.11 0.4634 DeepLPF[42] 18.44 0.7431
w/DiffLoss 14.66 0.4064 w/DiffLoss 17.67 0.4710 w/DiffLoss 19.51 0.7437
TaylorFormer*[47] 15.02 0.5178 EfDeRain*[22] 23.41 0.7524 IAT*[11] 19.89 0.7371
w/DiffLoss 15.19 0.5326 w/DiffLoss 24.54 0.7656 w/DiffLoss 20.11 0.7378
5.1 Experiment Setup
Datasets. We train and evaluate our models on both synthetic and real-world image restoration datasets, including low light enhancement, image deraining and image dehazing. For real-world challenging scenes, we adopt LOL dataset[61] for low light enhancement and Dense-Haze dataset[1] for image dehazing. For image deraining, we adopt Rain13K dataset[6] for training and Rain100H dataset[65] for testing.
Finally, we use CUB dataset[56] for image classification task, and the degraded CUB dataset is obtained through synthetic methods from [25] to simulate degraded conditions, such as fog, rain, and low-light scenarios.
Figure 6: Comparison of visual results on Rain100H and LOL datasets. (a):EfDeRain; (b): RCD-Net; (c): IAT; (d): DeepLPF. Please zoom in for best view.
Comparison Baseline Methods. We choose several classical and SOTA restoration methods as baselines, including IAT[11], DeepLPF[42] for low light enhancement, EfDeRain[22], RCD-Net[57] for image deraining, and TaylorFormer[47], FSDGN [67] for image dehazing. We separately train these baselines with and without our DiffLoss, with the same settings and implementation details. We compare these two settings on the above baselines qualitatively and quantitatively.
Figure 7: Comparison of visual results on Dense-Haze dataset. (e): FSDGN; (f): Taylorformer. Please zoom in for best view.
Implementation Details. We choose existing restoration networks listed above as backbone. For the DiffLoss, we adopt the unconditional diffusion model pre-trained on ImageNet dataset by [13]. Both the network architecture of restoration network and diffusion model need no modification. Besides, we fix the parameters of diffusion model. During the training stage, we use ADAM as the optimizers with the learning rate set to $1\times 10^{-4}$. The batch and patch sizes are set to $4$ and $256\times 256$, respectively. The parameter size of diffusion model is 552.81M. For image classification, we use VGG16[54] and Resnet50[24] pre-trained on clean CUB dataset as the recognition models to evaluate images restored by different methods. All the restoration models are trained with RTX 4090 GPU.
Evaluation Metrics. We evaluate our method on two different metrics: PSNR, and SSIM [59], which well-known image quality assessment indicators.
5.2 Comparison with Baseline Methods on Several Benchmarks.
First of all, need to note that our method improves the naturalness of restored results, instead of substantially removing more degradation.
Comparison on real-world degradation images. Tab. 1 compares the quantitative results of different methods on Dense-Haze, Rain100H and LOL datasets for image dehazing, image deraining and low light enhancement, respectively. On these datasets, Baselines with DiffLoss achieves better performance with most of metrics. The results on these challenging real-world datasets effectively demonstrate the advantages and effectiveness of our approach.
We also specifically utilized the efficient model, including TaylorFormer, EfDeRain and IAT, and the experiments revealed that our method enables the efficient model to achieve better image restoration. This is attributed to the inherent generative capability of the diffusion model, which helps the efficient model learn the distribution of natural images.
We also show the visual comparison with other method on real degradation images sampled from testing sets of Dense-Haze, Rain100H and LOL datasets. The visual results are presented in  Fig. 6 and Fig. 7. The complex degradation distribution makes these datasets extremely challenging, and existing methods on these datasets usually suffer from unnaturalness problems. As shown in  Fig. 6 and Fig. 7, these baselines produce relatively pleasing results, however, artifacts and blurs still emerge. In contrast, after being empowered by the DiffLoss, they produce results more visually pleasing than the baseline results. For better visualization, we denote the obvious region with the red rectangular box.
[…]
Config dehazing deraing low light enhancement
Top-1 V Top-1 R Top-1 V Top-1 R Top-1 V Top-1 R
w/o DiffLoss 37.5940 24.3966 68.6724 78.8966 15.2069 28.2069
w/ DiffLoss 46.7105 37.5906 70.0517 79.5690 35.7241 53.5690
Improvement on Image Classification. We have demonstrated in Fig. 2 that h-space possesses semantic attributes. Therefore, we utilize h-space as a loss to preserve the semantic information for low level tasks. We synthesized degraded images with low-light, haze, and rain using the origin CUB dataset. We train the model using common low-level datasets as mentioned before with and without DiffLoss, and use the trained model to restore the degraded CUB dataset. We choose to use Taylorformer, EfDeRain and IAT for different degradations. Finally, we employ a pre-trained VGG16[54] and Resnet50[24] networks to evaluate the accuracy of classification. From the Tab. 2, it is evident that with the assistance of our method, the model can preserve more semantic information after the restoration process.
Figure 8: Graph of PSNR with different weight of DiffLoss during training process on Dense-Haze [1] dataset with MSBDN [14] as baseline. We train the baseline from scratch and choose the best performance in the first 40K iterations.
Table 3: The results of ablated models on Dense-Haze [1] dataset with GridNet [38] as baseline.
[…]
a 429.73 14.010 0.3681
b 343.31 14.824 0.4429
c 338.63 14.743 0.4385
d (ours) 293.01 14.907 0.4666
5.3 Ablation study
In this section, we perform several ablation studies to analyze the effectiveness of the proposed DiffLoss on Dense-Haze dataset. The studies include the following ablated models: (a) With constraints on $\hat{x}_{0}$ only. (b) With constraints on $\hat{\hat{x}}_{0}$ only. (c) With constraints on $\hat{\varepsilon}$ and $\hat{x}_{t-1}$. (d) Ours (final setting). These models are trained using the same training setting as our method. The performance of these models is summarized in Table 3. Obviously, every design component in DiffLoss can elevate the performance, and constraining on $\hat{\varepsilon}$ and $\hat{x}_{t-1}$ performs similarly to constraining on $\hat{x}_{0}$ and $\hat{\hat{x}}_{0}$.
We also experiment with different loss value to get the optimal one. As shown in Fig. 8, we experiment with different weight of DiffLoss and plot the PSNR-Loss weight curve. Similar to perceptual and adversarial loss, DiffLoss gets the optimal performance with loss weight $\gamma\in[0.0005,0.005]$. In our method, we choose the loss weight of DiffLoss to be 0.001. This experiment is conducted with $\lambda=0$. The experiment of optimal choice for $\lambda$ and more comparison presented in the supplementary material.
6 Conclusion
In this paper, we propose a new perspective from the correlation of degraded and natural image distribution that achieves effective image restoration. To achieve this, inspired by the powerful capability of the diffusion model for natural image sampling and generation, we embed the pre-trained diffusion model into restoration network as a auxiliary training mechanism to empower the learning capability and semantic attributes of neural networks for effective naturalness image restoration.
By equipping existing restoration networks with the DiffLoss in the training stage, we can substantially elevate their performance and yield more natural and recognition-aware restored images without involving additional computations in the inference stage.